# Efficient Transformed Gaussian Process State-Space Models for Non-Stationary High-Dimensional Dynamical Systems

### v2

@@ I Introduction
State-space models (SSMs) offer a versatile framework for capturing dynamical systems where latent states evolve based on internal dynamics and external inputs [1]. For systems with well-defined dynamics, Bayesian filtering methods—including the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF)—enable sequential state estimation by integrating prior knowledge with new observations, thus facilitating reliable predictions in uncertain conditions. This capability makes SSMs highly applicable in fields such as robotics, finance, control systems, and signal processing, where modeling temporal dependencies is essential [2, 3, 4]. Yet, in many complex real-world scenarios, the lack of precise knowledge about system dynamics significantly limits traditional SSM methods, particularly for nonlinear, high-dimensional systems like climate models, physiological processes, robotics, and neural dynamics [5].
@@
To address this challenge, numerous data-driven modeling and system identification approaches have been proposed. Deterministic models, such as deep state-space models (DSSMs), employ neural networks to learn system dynamics and have gained popularity for their ability to capture high-dimensional, nonlinear behaviors [6]. However, these models often demand extensive training datasets to mitigate overfitting, exhibit limited generalization in uncertain environments, and lack interpretability in their predictions [7]. Consequently, despite their theoretical potential, DSSMs frequently underperform in practical applications [8].
@@
In contrast, stochastic models incorporating flexible random process priors offer a robust alternative within a Bayesian learning framework [9]. A prominent example is the class of Gaussian process state-space models (GPSSMs), which utilize Gaussian processes (GPs) to provide interpretable uncertainty quantification and inherent regularization, making them extraordinary for modeling safety-critical dynamical systems [10, 11]. Consequently, GPSSMs have been extensively utilized in various applications, including human pose and motion learning [12], robotics and control learning [13], reinforcement learning [14, 15], target tracking and navigation [16], and magnetic-field sensing [17].
@@
Research on GPSSMs has also focused on advancing learning and inference methodologies. Early works either assumed a pre-learned GPSSM and concentrated solely on latent state inference—an impractical assumption for many real-world applications [18, 19, 20]—or employed maximum a posteriori (MAP) estimation to jointly infer the model and the latent states [12, 21], which is prone to overfitting. To address these issues, a pivotal development came with the fully Bayesian treatment of GPSSMs using particle Markov chain Monte Carlo (PMCMC) methods [22], though the inherent cubic computational complexity of GP remains prohibitive for long time series. Subsequent efforts [23, 24, 25, 26] adopted reduced-rank GP approximations to alleviate the computational demands, yet PMCMC scalability remains a challenge. This has led to a growing interest in variational inference methods [10, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39], which leverage sparse GP approximations with inducing points to enhance scalability [40, 41].
@@
However, existing GPSSMs, which employ multiple independent GPs to model system dynamics, remain limited to low-dimensional state spaces and face fundamental challenges when scaling to high-dimensional systems. Specifically, in variational GPSSMs, the inherent cubic computational complexity of the GPs scales linearly with the state dimension, while the number of parameters grows quadratically. This computational and parametric complexity seriously hinders the applicability of GPSSMs to high-dimensional systems. Additionally, all GPSSMs inherit a fundamental limitation, namely the use of stationary random process priors for modeling system dynamics. Such stationary assumptions are often inadequate for capturing time-varying and complex transition functions, ultimately limiting GPSSMs’ ability to model complex dynamical systems accurately.
@@
To address these challenges, we propose a novel approach for learning and inference in GPSSMs tailored to non-stationary, high-dimensional dynamical systems. Our primary contribution is the development and integration of an efficient yet flexible implicit process prior [42] for the state transition function. This prior combines the strengths of GPs, normalizing flows [43], and (Bayesian) neural networks, significantly enhancing modeling capabilities in high-dimensional state spaces. Our key contributions are summarized as follows:
@@
  • Instead of employing multiple independent GPs for the transition function, we propose applying an efficient transformed Gaussian process (ETGP), an implicit random process that combines a shared GP with normalizing flows. Crucially, the flow parameters are made input-dependent through neural network modeling. This novel modeling approach offers two key advantages: (1) it captures complex, non-stationary transition functions that standard GP priors cannot represent, and (2) by utilizing only a single shared GP, it mitigates the computational and parametric complexities of traditional GPSSMs in high-dimensional spaces, enabling effective scaling to higher dimensions.
@@
  • For efficient inference, we develop a scalable variational algorithm to jointly approximate the posterior distributions of both the ETGP and the latent states. Unlike standard GPs, the implicit nature of the ETGP prevents direct parametrization of the function-space posterior. To address this challenge, we explicitly approximate the posterior over the underlying GP and the associated neural network parameters, which together define the ETGP posterior. Additionally, we incorporate the EnKF into the variational inference framework, eliminating the need to parameterize the variational distributions over latent states. This hybrid formulation thus not only offers accurate state estimation but also substantially reduces the computational cost of the parameter optimization.
@@
The remainder of this paper is structured as follows. Section II introduces the preliminaries of state-space models and outlines key challenges in existing GPSSMs. Section III presents our proposed efficient GPSSM framework for high-dimensional dynamical systems, including the associated variational inference algorithm. Experimental results are detailed in Section IV, followed by concluding remarks in Section V. Technical proofs and derivations are deferred to the Appendix.
@@ II-A State-Space Models (SSMs)
A generic SSM describes the probabilistic relationship between the latent state ${\mathbf{x}}_{t}\in\mathbb{R}^{d_{x}}$ and the observation ${\mathbf{y}}_{t}\in\mathbb{R}^{d_{y}}$. It can be mathematically expressed as:
@@
Here, the latent states follow a Markov process, meaning that at any time step $t\in\mathbb{N}$, the next state ${\mathbf{x}}_{t+1}$ depends only on the current state ${\mathbf{x}}_{t}$ and the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$. In this paper, the emission function is assumed to be linear, with the coefficient matrix $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$. For systems with nonlinear emission functions, the latent state can be augmented to a higher dimension, transforming the system into one with a linear emission function [30], which mitigates the non-identifiability issues commonly encountered in data-driven SSMs [10]. Both the latent states and observations are assumed to be perturbed by zero-mean Gaussian noise, with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
@@ II-B1 Gaussian process (GP)
The GP is a generalization of the Gaussian distribution over infinite index sets, enabling the specification of distributions over functions $\tilde{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$. A GP is fully characterized by its mean function, often set to zero, and its covariance (kernel) function $k({\mathbf{x}},{\mathbf{x}}^{\prime})$, which includes a set of hyperparameters $\bm{\theta}_{gp}$ that need to be optimized for model selection [9]. According to the definition of a GP, the function values $\tilde{{\mathbf{f}}}=\{\tilde{f}({\mathbf{x}}_{i})\}_{i=1}^{N}\in\mathbb{R}^{N}$ at any finite set of points $\mathbf{X}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{x}}$ follow a joint Gaussian distribution, i.e.,
@@
where $\mathbf{K}\in\mathbb{R}^{N\times N}$ is the covariance matrix evaluated on the finite input $\mathbf{X}$ with $[\mathbf{K}]_{i,j}\!=\!k({{{\mathbf{x}}}}_{i},{{\mathbf{x}}}_{j})$.
Given the observed function values $\tilde{{\mathbf{f}}}$ at the input $\mathbf{X}$, the GP prediction distribution, $p(\tilde{f}({\mathbf{x}}_{*})|\tilde{{\mathbf{f}}},\mathbf{X})$, at any new input ${{\mathbf{x}}}_{*}$, is also Gaussian, and fully characterized by the posterior mean $\xi({{\mathbf{x}}}_{*})$ and the posterior variance $\Xi({{\mathbf{x}}}_{*})$. Concretely,
@@
where $\mathbf{K}_{\mathbf{x}_{*},\mathbf{X}}$ is the cross covariance matrix evaluated on the new input ${{\mathbf{x}}}_{*}$ and the observed input ${\mathbf{X}}$.
@@ II-B2 Gaussian process state-space model
Placing a GP prior over the transition function $\bm{f}(\cdot)$ in the classic SSM (see Eq. (1)) leads to the following salient GPSSM [10]:
@@
where the initial state prior distribution $p({\mathbf{x}}_{0})$ is assumed to be known and follows a Gaussian distribution. The graphical representation is depicted in Fig. 1.
In GPSSMs, when the state dimension $d_{x}>1$, the transition $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$ is typically modeled using $d_{x}$ mutually independent GPs. Each independent GP represents a dimension-specific function ${f}_{d}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$, and the multivariate output is expressed as
@@
where each independent GP is characterized by a distinct kernel function and associated hyperparameters [30]. The main challenge in GPSSMs is the simultaneous learning of the transition function and noise parameters, i.e., learning $[{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}]$, while inferring both the GPs and the latent states of interest.
@@ II-C Challenges in Existing Variational GPSSMs
Note that in Eq. (3), evaluating the GP posterior involves inverting the covariance matrix $\mathbf{K}$, resulting in a computational cost of $\mathcal{O}(N^{3})$, which becomes prohibitive for large $N$. To address this issue, variational GPSSMs typically adopt sparse GP approximations [40, 41], a ubiquitous technique in GP-based models [44, 10] due to its analytical and computational tractability. The general idea of the sparse GP involves augmenting the corresponding GP by introducing a set of inducing points $\mathbf{Z}\!=\!\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ and ${\mathbf{u}}\!=\!\{\mathrm{u}_{i}\}_{i=1}^{M},M\ll N,$ that will serve as a surrogate for the GP. Here, ${\mathbf{u}}\!\sim\!{\cal N}(\bm{0},\mathbf{K}_{\mathbf{Z},\mathbf{Z}})$ and ${\mathbf{z}}_{i}\!\in\!\mathbb{R}^{d_{x}}$. Specifically, the augmented GP prior is given by [^1] With slight notation abuse, we denote the distribution of GP prior as $p(\tilde{f})$.,
@@
We can approximate the corresponding GP posterior by explicitly specifying a free Gaussian density on the function values $\mathbf{u}$ at $M$ locations, $q(\mathbf{u})\!=\!\mathcal{N}\left(\mathbf{u}|\mathbf{m},\mathbf{S}\right)$, and using the prior conditional for the rest of the GP, i.e.,
@@
This leads to an explicit form for the approximate GP posterior
@@
resulting in a computational complexity that scales as $\mathcal{O}(M^{3})$ instead of $\mathcal{O}(N^{3})$.
Nevertheless, existing variational GPSSMs still face significant computational and modeling challenges, which are listed below.
  • Modeling multi-dimensional state transition function using multiple separate GP results in a computational complexity of $\mathcal{O}(d_{x}M^{3})$ in GPSSMs, where the complexity grows linearly with the latent state dimension $d_{x}$.
  • The variational approximation of the GP posterior introduces a number of variational parameters that scale quadratically with the state dimensionality $d_{x}$. Specifically, the variational parameters in multi-dimensional GPSSMs include inducing inputs $\{\mathbf{Z}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times d_{x})}$, variational means $\{\mathbf{m}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times M}$, and covariance matrices $\{\mathbf{S}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times M)}$ for the $d_{x}$-dimensional inducing outputs $\{\mathbf{u}_{d}\}_{d=1}^{d_{x}}$. As the state dimension increases, the quadratic scaling of inducing inputs can lead to computational intractability, while the scaling of the variational means and covariance matrices further compounds the computational burden [38].
  • Existing GPSSMs commonly place a stationary GP prior over the transition function, as shown in Eq. (4a), to model a time-invariant transition function. However, this approach may fail to adequately capture non-stationary dynamics, leading to inaccuracies in scenarios where the transition function evolves over time. Moreover, the use of independent GPs for each dimension of the transition function overlooks dependencies between outputs.

### v3

@@ I Introduction
State-space models (SSMs) provide a flexible framework for modeling dynamical systems, where latent states evolve based on internal dynamics and external inputs [1]. When system dynamics are fully known, Bayesian filtering methods—such as the Kalman filter (KF), extended Kalman filter (EKF), ensemble Kalman filter (EnKF), and particle filter (PF)—enable sequential state estimation by combining prior knowledge with incoming observations, yielding reliable predictions under uncertainty. This capability makes SSMs widely applicable in robotics, finance, and signal processing, where modeling temporal dependencies is crucial [2, 3, 4].
However, in complex real-world systems, including nonlinear and high-dimensional problems such as climate modeling, physiological processes, and neural dynamics, the underlying dynamics may be too intricate or poorly understood to specify a physics-informed or explicitly parameterized model. In these cases, traditional filtering methods cannot be applied in their standard form and must instead rely on data-driven approaches to learn the dynamics from observations, complicating the task of accurate and robust state estimation [5].
@@
To address this challenge, a variety of data-driven system identification approaches have been proposed. Deterministic models, such as deep state-space models (DSSMs), employ neural networks to learn system dynamics and have gained popularity for their ability to capture high-dimensional, nonlinear behaviors [6]. However, these models often require large training datasets to avoid overfitting, and their generalization performance may degrade in uncertain or data-scarce environments. Moreover, their predictions can be difficult to interpret, especially when compared to more structured probabilistic models [7]. As a result, DSSMs may face challenges in practical deployment despite their modeling capacity [8].
@@
Stochastic models based on flexible random process priors provide a principled alternative within a Bayesian learning framework [9]. A prominent example is the class of Gaussian process state-space models (GPSSMs), which leverage Gaussian processes (GPs) to capture uncertainty and provide implicit regularization through prior structure. These properties make GPSSMs attractive for applications where uncertainty quantification and data efficiency are critical [10, 11]. GPSSMs have been applied in diverse domains, including human pose and motion learning [12], robotics and control [13], reinforcement learning [14, 15], target tracking and navigation [16], and magnetic-field sensing [17].
@@
Research on GPSSMs has also focused on advancing learning and inference methodologies. Early works either assumed a pre-learned GPSSM and concentrated solely on latent state inference—an impractical assumption for many real-world applications [18, 19, 20]—or employed maximum a posteriori (MAP) estimation to jointly infer the model and the latent states [12, 21], which is prone to overfitting. To address these issues, a important development came with the fully Bayesian treatment of GPSSMs using particle Markov chain Monte Carlo (PMCMC) methods [22], though the inherent cubic computational complexity of GP remains prohibitive for long time series. Subsequent efforts [23, 24, 25, 26] adopted reduced-rank GP approximations to alleviate the computational demands, yet PMCMC scalability remains a challenge. This has led to a growing interest in variational inference methods [10, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] that can leverage sparse GP approximations with inducing points to enhance scalability [40, 41].
@@
However, existing GPSSMs, which employ multiple independent GPs to model system dynamics, remain limited to low-dimensional state spaces and face fundamental challenges when scaling to high-dimensional systems. Specifically, in variational GPSSMs, each output dimension of the state transition function is modeled by a separate GP, leading to a total computational cost that scales linearly with the state dimensionality. Since each GP has a cubic complexity in the number of data points due to covariance matrix inversion, the overall computational burden becomes prohibitive. Moreover, the number of model parameters grows quadratically with the state dimension, further exacerbating scalability issues. Additionally, all GPSSMs rely on stationary GP priors, which assume that the statistical properties of the transition function are invariant under input translation. This assumption often fails for systems with time-varying or non-stationary dynamics, limiting the ability of GPSSMs to accurately capture complex real-world behaviors.
@@
To address these challenges, we propose a novel approach for learning and inference in GPSSMs tailored to non-stationary, high-dimensional dynamical systems. Our primary contribution is the development and integration of an efficient yet flexible implicit process prior [42, 43] for the state transition function. This prior combines the strengths of GPs, normalizing flows [44], and (Bayesian) neural networks, significantly enhancing modeling capabilities in high-dimensional state spaces. Our key contributions are summarized as follows:
@@
  • Instead of employing multiple independent GPs for the transition function, we propose applying an efficient transformed Gaussian process (ETGP) [42], an implicit random process that combines a shared GP with normalizing flows. Crucially, the flow parameters are made input-dependent through neural network modeling. This novel modeling approach offers two key advantages: (1) it captures complex, non-stationary transition functions that standard GP priors cannot represent, and (2) by utilizing only a single shared GP, it mitigates the computational and parametric complexities of traditional GPSSMs in high-dimensional spaces, enabling effective scaling to higher dimensions.
@@
  • For efficient inference, we develop a scalable variational algorithm to jointly approximate the posterior distributions of both the ETGP and the latent states. Unlike standard GPs, the implicit nature of the ETGP prevents direct parametrization of the function-space posterior. To address this challenge, we explicitly approximate the posterior over the underlying GP and the associated neural network parameters, which together define the ETGP posterior. Additionally, we derive a generic and assumption-free variational lower bound that can incorporate an EnKF into the variational inference framework, eliminating the need to parameterize the variational distributions over latent states. This hybrid formulation thus not only offers accurate state estimation but also substantially reduces the computational cost of the parameter optimization.
@@
The remainder of this paper is structured as follows. Section II introduces the preliminaries of state-space models and outlines key challenges in existing GPSSMs. Section III presents our proposed efficient GPSSM framework for high-dimensional dynamical systems, including the associated variational inference algorithm. Experimental results are detailed in Section IV, followed by concluding remarks in Section V. Supporting results and proofs are deferred to the Appendix and supplementary materials.
@@ II-A State-Space Models (SSMs)
An SSM defines a probabilistic dynamical system by specifying how latent states ${\mathbf{x}}_{t}\in\mathbb{R}^{d_{x}}$ evolve over time and generate observations ${\mathbf{y}}_{t}\in\mathbb{R}^{d_{y}}$. It is commonly expressed as:
@@
The latent states evolve according to a first-order Markov process, meaning that at any time step $t\in\mathbb{N}$, the next state ${\mathbf{x}}_{t+1}$ depends only on the current state ${\mathbf{x}}_{t}$ through the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}^{d_{x}}$. In this work, the emission function is assumed to be linear, parameterized by a matrix $\bm{C}\in\mathbb{R}^{d_{y}\times d_{x}}$. For systems with nonlinear emissions, the latent state can be augmented to a higher-dimensional representation, effectively rendering the emission function linear [30]. This reparameterization helps mitigate non-identifiability issues frequently encountered in data-driven SSMs [10]. Both latent states and observations are perturbed by zero-mean Gaussian noise, with covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, respectively.
@@ II-B1 Gaussian process (GP)
The GP is a generalization of the Gaussian distribution over infinite index sets, enabling the specification of distributions over functions $\tilde{f}(\cdot):\mathbb{R}^{d_{x}}\mapsto\mathbb{R}$. A GP is fully characterized by its mean function, often set to zero, and its covariance (kernel) function $k({\mathbf{x}},{\mathbf{x}}^{\prime})$, which includes a set of hyperparameters $\bm{\theta}_{gp}$ optimized during model selection [9]. By definition, the function values $\tilde{{\mathbf{f}}}=\{\tilde{f}({\mathbf{x}}_{i})\}_{i=1}^{N}\in\mathbb{R}^{N}$, evaluated at a finite set of inputs $\mathbf{X}=\{{\mathbf{x}}_{i}\}_{i=1}^{N}\in\mathbb{R}^{N\times d_{x}}$, follow a joint Gaussian distribution:
@@
where $\mathbf{K}\in\mathbb{R}^{N\times N}$ is the covariance matrix with entries $[\mathbf{K}]_{i,j}=k({\mathbf{x}}_{i},{\mathbf{x}}_{j})$.
Given observed function values $\tilde{{\mathbf{f}}}$ at inputs $\mathbf{X}$, the GP prediction at a new input ${\mathbf{x}}_{*}$, denoted $p(\tilde{f}({\mathbf{x}}_{*})\mid\tilde{{\mathbf{f}}},\mathbf{X})$, is also Gaussian, with posterior mean $\xi({\mathbf{x}}_{*})$ and posterior variance $\Xi({\mathbf{x}}_{*})$ given by:
@@
where $\mathbf{K}_{{\mathbf{x}}_{*},\mathbf{X}}$ is the cross-covariance vector between the test input ${\mathbf{x}}_{*}$ and the training inputs $\mathbf{X}$.
@@ II-B2 Gaussian process state-space model
Placing a GP prior over the transition function $\bm{f}(\cdot)$ in the classic SSM (see Eq. (1)) yields the GPSSM [10]:
@@
where the initial state $p({\mathbf{x}}_{0})$ is assumed known and Gaussian. A graphical representation is shown in Fig. 1.
When the state dimension $d_{x}>1$, the transition function $\bm{f}(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{x}}$ is typically modeled using $d_{x}$ independent GPs. Each component function $f_{d}(\cdot):\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}$ maps to one output dimension:
@@
where each GP has its own kernel function and hyperparameters [30]. The central challenge in GPSSMs lies in jointly learning the transition dynamics and noise parameters, i.e., learning $[{{\bm{\theta}}}_{gp},\mathbf{Q},\mathbf{R}]$, while simultaneously inferring both the latent states and the underlying functions.
@@ II-C Challenges in Existing Variational GPSSMs
Note that in Eq. (3), evaluating the GP posterior requires inverting the covariance matrix $\mathbf{K}$, incurring a computational cost of $\mathcal{O}(N^{3})$, which becomes prohibitive for large $N$. To address this, variational GPSSMs typically employ sparse GP approximations [40, 41], a widely used technique in GP-based models [45, 10] due to its analytical tractability and reduced computational cost. The core idea is to augment the GP with a smaller set of inducing points, which act as a sparse approximation of the full GP.
More specifically, we can introduce $M\!\ll\!N$ inducing inputs $\mathbf{Z}=\{{\mathbf{z}}_{i}\}_{i=1}^{M}$ (where ${\mathbf{z}}_{i}\in\mathbb{R}^{d_{x}}$) and their corresponding inducing outputs ${\mathbf{u}}=\{\mathrm{u}_{i}\}_{i=1}^{M}=\{\tilde{f}({\mathbf{z}}_{i})\}_{i=1}^{M}$, which inherit the same GP prior as the original process, i.e., ${\mathbf{u}}\sim\mathcal{N}(\bm{0},\mathbf{K}_{\mathbf{Z},\mathbf{Z}})$. The augmented GP prior can be written as: [^1] With slight notation abuse, we denote the distribution of GP prior as $p(\tilde{f})$.
@@
To approximate the posterior, a variational distribution is defined over the $M$ inducing outputs, $q({\mathbf{u}})=\mathcal{N}({\mathbf{u}}\mid\mathbf{m},\mathbf{S})$, while the prior conditional is used for the rest of the GP:
@@
This leads to an explicit expression for the variational approximation to the GP posterior:
@@
reducing the computational complexity from $\mathcal{O}(N^{3})$ to $\mathcal{O}(M^{3})$. Nevertheless, existing variational GPSSMs face several computational and modeling challenges, outlined below:
  • Modeling multi-dimensional state transition function using separate GPs for each output dimension results in a computational complexity of $\mathcal{O}(d_{x}M^{3})$, scaling linearly with the latent state dimension $d_{x}$.
  • The variational approximation introduces a number of variational parameters that scale quadratically with the state dimensionality $d_{x}$. Specifically, variational GPSSMs define inducing inputs $\{\mathbf{Z}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times d_{x})}$, variational means $\{\mathbf{m}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times M}$, and covariance matrices $\{\mathbf{S}_{d}\}_{d=1}^{d_{x}}\in\mathbb{R}^{d_{x}\times(M\times M)}$ for the $d_{x}$-dimensional inducing outputs $\{\mathbf{u}_{d}\}_{d=1}^{d_{x}}$. As the state dimension increases, the quadratic scaling of inducing inputs and the associated variational parameters leads to significant computational overhead [38].
  • Most GPSSMs impose a stationary GP prior over the transition function, as shown in Eq. (4a), to model time-invariant dynamics. However, such stationary assumptions may fail to capture time-varying or context-dependent behavior [^2] Here, “stationarity” refers to the GP prior over the transition function, not the observed time series. A GPSSM with a stationary kernel may still model non-stationary observed data.. Additionally, modeling each output dimension with an independent GP ignores correlations across dimensions, potentially limiting modeling capacity.